DIRECTIONS: Ask for the probability of a randomly selected item being less than a value that is larger than your mean(please show me the steps in how this is solved )

I sampled 8 books and found the mean length (# of pages) to be 276 with a sample standard deviation of 34 pages. Find the probability that a randomly selected book has less than 150 pages.
Answer #1

Solution :

Given that ,

mean = = 276

standard deviation = s = 34

n = 8

degrres of freedom = n - 1 = 8 - 1 = 7

P(x < 150) = P((x - ) / s < (150 - 276) / 34)

= P(t < -3.71)

Using t table

= 0.0038